OpenStudy (ammarah):

Factor the polynomial completely using any method. Sum or difference of cubes or by grouping. x^3-512 -3y^3-24 y^4-81

OpenStudy (ammarah):

@nikato

OpenStudy (nikato):

i'll try. kinda forgot about this from last year.. but do you know like the special frmulas for factoring cubes

OpenStudy (anonymous):

\[(x-8)(x ^{2} + 8x + 64)\]

OpenStudy (nikato):

a3 + b3 = (a + b)(a2 – ab + b2) a3 – b3 = (a – b)(a2 + ab + b2) do you know this? special formulas for cubes

OpenStudy (anonymous):

On the second one if you factor out -3 it will help a lot.

OpenStudy (ammarah):

so the second one is -3(y^3+8)

OpenStudy (anonymous):

Yeah. Where a = y and b = 2

OpenStudy (ammarah):

so what would that be?

OpenStudy (anonymous):

(y + 2)(y^2 + 2y + 4)

OpenStudy (ammarah):

how about the thir d one

OpenStudy (ammarah):

wait ill tell u what i got

OpenStudy (anonymous):

Not sure about that one

OpenStudy (nikato):

third one... its a special kind of factoring using this fromula (x-y)(x+y)=x^2-y^2

OpenStudy (ammarah):

idk how t solve it...

OpenStudy (anonymous):

\[(y ^{2} + 9)(y ^{2} -9)\] and solve for that

OpenStudy (ammarah):

how do i solve?

OpenStudy (anonymous):

So try to factor each one individually.

OpenStudy (anonymous):

First one doesn't factor.

OpenStudy (ammarah):

y+3 and y-3

OpenStudy (anonymous):

\[(y ^{2} + 9)(y + 3)(y -3)\]

OpenStudy (ammarah):

yes

OpenStudy (anonymous):

Good job

OpenStudy (ammarah):

i need more help on a diff question

OpenStudy (anonymous):

sure what is it